A car accelerates from zero to a speed of 36 km/h in 15 s.

i.

Calculate the acceleration of the car in m/s2.

ii.

If the acceleration is assumed to be constant, how far will the car travel in 1 minute ?

iii.

Calculate the speed of the car after 1 minute.


1
Expert's answer
2022-06-13T08:23:19-0400

Given:

"v_i=0\:\rm m/s"

"v_f=36\:\rm km/h=10\:\rm m/s"

"t=15\:\rm s"

"t_1=60\:\rm s"


(i)

"a=\frac{v_f-v_i}{t}=\frac{10-0}{15}=0.67\:\rm m/s^2"

(ii)

"d=v_it_1+at_1^2/2\\
=0*60+0.67*60^2/2=1200\:\rm m"

(iii)

"v=v_it_1+at_1\\
=0*60+0.67*60=40\:\rm m/s"